media%2Fb34%2Fb34b57ad-a0be-4d59-874d-9d

media%2F152%2F152e240b-9281-42fd-b167-71
Answer #1

1)Because the pressure is the highest at the bottom and probably the thickness of the levee was independent of depth.

2)Mass=rho*pi*r^2*h=7870*pi*0.01425^2*0.858=4.31 kg.

Clearly a cart is not needed.

3)Density of sun=mass/volume=1.99*10^30/(4.3*pi*(6.96*10^8)^3)=436.93kg/m^3

4)Gauge pressure=rho*g*h=1000*9.81*3200Pa=3.139*10^7Pa=309.8 atm

5)Gauge pressure at interface=rho1*g*h1=600*9.81*0.12=706.3Pa

Gauge pressure at bottom=rho1*g*h1+rho2*g*h2=600*9.81*0.12+1000*9.81*0.25=3158.8Pa

6)Pressure=F/A=1.5/(pi*0.0041^2)=28404Pa

h=P/rho*g=28404/(1000*9.81)=2.90m

7)Buoyancy=17.5-11.2=6.3N

Volume of water=buoyant force/(rho*g)=6.3/(1000*9.8)=6.43*10^-4m^3

Density=Mass/volume=(weight/g)/(volume)=(17.5/9.81*6.43*10^-4)=2774kg/m^3

8)Buoyant force=39.2-28.4=10.8N

Volume=10.8/(9.81*1000)=1.101*10^-3m^3

Buoyant force in other liquid=39.2-18.6=20.6N

Density=buoyant force/(g*volume)=20.6/(9.81*1.101*10^-3)=1907kg/m^3